40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Nov 26 07:22:20 2011, suresh, Update, Computers,  
    Reply  Mon Nov 28 06:53:30 2011, kiwamu, Update, CDS, c1sus shutdown 
       Reply  Mon Nov 28 16:46:55 2011, kiwamu, Update, CDS, c1sus is now up 
Message ID: 6020     Entry time: Mon Nov 28 06:53:30 2011     In reply to: 6016     Reply to this: 6026
Author: kiwamu 
Type: Update 
Category: CDS 
Subject: c1sus shutdown 

I have restarted the c1sus machine around 9:00 PM yesterday and then shut it down around 4:00 AM this morning after a little bit of taking care of the interferomter.

Quote from #6016

c1sus has been shutdown so that the optics dont bang around.  This is because the watch dogs are not working.

ELOG V3.1.3-